OpenStudy (anonymous):

Solve: x^4 + 4x^3 - x - 4 = 0

OpenStudy (anonymous):

x^3(x+4)-1(x+4)=0 (x^3-1)(x+3)=0 so x=1,-3

OpenStudy (anonymous):

the real zeros are 4 and -1 (from the graph of it), plus there are two complex zeros. :/

OpenStudy (anonymous):

Yeap, I havenot considered complex numbers here.........

OpenStudy (anonymous):

sorry (x+4)

OpenStudy (anonymous):

so x=-4

OpenStudy (anonymous):

Oh, right. But I still need the complex zeros...

OpenStudy (anonymous):

-0.5+0.87i -0.5-0.87i

OpenStudy (anonymous):

Might you have the exact number/fraction for that .87?